OpenStudy (anonymous):

Which is an equation of a circle with center (2, 7) and radius 4? A. (x + 2)^2 + (y + 7)^2 = 4 B. (x - 7)^2 + (y - 2)^2 = 16 C. (x - 2)^2 + (y - 7)^2 = 4 D. (x – 2)^2 + (y - 7)^2 = 16

OpenStudy (anonymous):

D.\[\left( x-h \right)^{2}+\left( y-k \right)^{2}=r ^{2}\] with centre (h,k) and radius=r

OpenStudy (anonymous):

I agree

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

yw